Writing 0,425 as an ordinary fraction in its simpliest form gives?


1
Expert's answer
2021-05-07T11:39:02-0400

Solution. To convert decimal to fraction


"0.425=\frac{425}{1000}=\frac{17}{40}"

Answer.


"0.425=\frac{17}{40}"